# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v3

@@ Appendix B Research Background
B.1 Any-Shot Cross-Task Generalization
Learning from zero-shot or few-shot examples has been identified as a crucial adaptation capability of the machine learner nowadays 70, 71. This work refers to traditional finetuning as many-shot learning and treats the individual example as each task in the MPTS setup. As finetuning techniques have been widely discussed in the field 61, we skip this part in the background introduction.
@@ B.1 Any-Shot Cross-Task Generalization
This assesses the machine learner’s generalization capability when directly deploying in unseen scenarios without the help of a support dataset. Such a cross-task generalization is commonly studied in computer vision 72, and the core of the relevant methods is effective semantic representation either from embedding-based methods 73, 74, 75 or generative-based methods 76, 77. In the era of the foundation models, the pretraining mechanism between multimodality also sometimes empowers the machine learner, such as CLIP 1, with zero-shot capability. When it comes to sequential decision-making, a commonly used technique is called DR 43, 9. In detail, DR places a distribution over environments for the agent to interact, deriving a generalist policy to play quickly in unseen scenarios.
@@
This examines the machine learner’s capability of resolving unseen tasks from some annotated examples as hints. Meta-learning, as the typical learning paradigm, has gained popularity over the past decade. It achieves few-shot adaptation by leveraging past experience and distilling knowledge to unseen but similar scenarios in a few-shot way 78. In brief, we categorize commonly seen methods into context-based, optimization-based, geometric-based, and others. (i) Formulated in an encoder-decoder structure, the context-based method resembles variational autoencoders and encodes the few-shot information into latent variables or embeddings. Typical ones are neural process families 55, 79, 80, 81, which aim to constitute exchangeable deep stochastic processes with neural networks. (ii) The optimization-based methods, with their versatile nature and ability to enable cross-task skill transfer, have piqued the interest and engagement of researchers in the field. For example, MAML 28, 82, 83, 84 reduces meta-learning to a bi-level optimization in the parameter space, and its extensions have been widely investigated in the field. (iii) The deep metric-based methods 85, 86 attempt to embed tasks into the latent space and are more suitable for few-shot image classification tasks. Besides, there are other families, such as hyper-networks 87, 88, recurrent meta-learning 89, etc.
@@ B.2 Dataset Curation and Task-Level Robustness
The risk minimization principle and the adopted sampling criteria are mainly entangled in optimization. Previously, Section 2.2 summarizes the commonly-used risk minimization principles in task episodic learning. Regardless of the adaptation concept and type, we will overview the existing task sampling strategies according to diverse views.
Data Curation in Typical Methods.
In adaptation learning literature, several task curation techniques are devised to pick up high-quality and difficult yet learnable examples 17, 29, 19, 30, 20, and empirical findings examine their role in yielding robustness enhancement and efficiency gains. There are other coreset methods that find a small portion of the task dataset to represent the full batch example’s utility, e.g., via gradient approximation in optimization 64, 18, 65. The studied framework brings a subproblem to solve, and its primary goal is to achieve data efficiency. Hence, the acquisition strategy used in MPTS can be viewed as the episodic coreset selection for robustness purposes.
@@
The CVaR or expected shortfall 31 is a statistical measure to assess the proportional worst-case performance of some models at certain levels. This is widely adopted in risk-averse applications. As implied in Definition 1, $\text{CVaR}_{\alpha}$ describes the expected risk under the normalized $(1-\alpha)$ proportional tail risk task distribution, and this work specifies the distribution in the task space. Meanwhile, the normalized tail task distribution $p_{\alpha}(\tau;\bm{\theta})$ can be viewed as a shifted result from the initial task distribution $p(\tau)$; hence, such a measure provides robustness quantification in the presence of task distribution shifts 16, 46, 33.
Definition 1 (Conditional Value-at-Risk (CVaR) 31)
[…]
Another strategy to evaluate the machine learner’s robustness is the performance in OOD tasks. This refers to the case when the training and the testing task distributions are different. Particularly, in DR and prompt-tuning scenarios, we also use the OOD tasks that never appear in the training task distribution to test the trained policy, and this setup corresponds to domain generalization, a type of substantial distribution shift 11.
B.3 Baseline Details
GDRM 35.
As implied in the main paper, this can be interpreted as a min-max optimization problem. The basic idea of increasing the machine learner’s robustness lies in assigning more probability mass to worst cases in a reweighted manner. That means in each iteration with the best selected $p_{\hat{g}}(\tau)$, the optimization problem is reduced as
@@ B.3 Baseline Details
Given a fixed number of tasks, GDRM will heuristically or dynamically group them into clusters and then perform a reweighting mechanism according to the evaluated risk. In task episodic learning, there is no task grouping operation as the task batch is reset after each iteration 16, 46. And the default computation of task-specific weigths is $\omega(\tau_{i})=\frac{\exp(\eta\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}))}{\sum_{b=1}^{\mathcal{B}}\exp(\eta\ell(\mathcal{D}_{\tau_{b}}^{Q},\mathcal{D}_{\tau_{b}}^{S};\bm{\theta}))}$, where $\eta$ is the temperature parameter and $\{\bm{\tau}_{b}\}_{b=1}^{\mathcal{B}}$ is the identifier of the task batch. The implementation detail can be found in https://github.com/kohpangwei/group_DRO.
DRM 16, 46.
Similar to GDRM, the optimization objective $\mathbb{E}_{p_{\alpha}(\tau;\bm{\theta})}\Big{[}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\Big{]}$ also defines the task distribution with constraints. As $p_{\alpha}(\tau;\bm{\theta})$ is the machine learner’ parameter $\bm{\theta}$dependent, a commonly used practical implementation is to evaluate the performance of the machine learner, rank the task-specific adaptation risk value, and then select $(1-\alpha)$ proportional worst cases to optimize.
In this work, we retain the setup in 16 and pick up the $-\mathcal{B}$ in optimization. This implies that the actual task batch to evaluate is $\frac{\mathcal{B}}{1-\alpha}$. And for fair comparison with MPTS and light computations, we retain the Monte Carlo estimator for the risk quantile in implementation, which is same as 16. To ensure stable training, for all benchmarks, we keep the actual task batch ^$\hat{\mathcal{B}}=2\mathcal{B}$ to evaluate and discard the easiest half before the machine learner’s optimization.
As revealed in works 90, 38, 16, the heuristic operation as the Evaluate-Rank-Filter or reweighting mechanism in GDRM is widely adopted for approximate optimization. For example, in task robust meta-learning scenarios, the prerequisite step in DR-MAML 16 is to execute gradient updates in the inner loop for all candidate tasks and then screen the tail task subset to meta-optimize according to the evaluation results.
Appendix C Task Formulations and Identifiers
@@ C.1 Tasks with Explicit Identifiers
In $K$-shot sinusoid regression 28, meta learners aim at quickly adapting the model to an unseen function $f(x)=a\sin(x-b)$ with the help of $K$ data points randomly sampled from the function. This case treats the amplitude and phase variables $(a,b)$ as the task identifier to configure the task. And the task distribution is induced by the uniform distribution over the task identifier.
@@
DR is a promising paradigm to achieve zero-shot adaptation in unseen scenarios, which is widely adopted in robotics 9 and computer vision 43. The basic idea is to train the machine learner in a wide variety of environments or deployment scenarios and then directly apply the learned model to new ones.
@@ C.2 Tasks with Implicit Identifiers
Nevertheless, in several scenarios, it is intractable to access the explicit task identifier. For example, in few-shot image classification, the task information is just the coupled support and query dataset $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$. Similarly, in PEFT for LLMs, the task can be in the form of the QA pair $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{Q}$. There is no explicit representation method, such as $\bm{\tau}$, for these tasks, which brings difficulty in building up the risk learner. Retaining the prior notation, the episodic task batch can be written as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$, where $\bm{\tau}$ of our interest is unobservable. Some experiments in this work, such as few-shot image classification and PEFT, encounter such embarrassment.
@@ Appendix D Auto-Encoding Adaptation Risk through Streaming VI
Note that the basis of MPTS is to establish the bridge between the task identifier and the adaptation risk value over the course of the machine learner’s optimization. This implies we are seeking a lightweight stochastic risk function in Definition 2 to approximate the posterior $p(\ell|\bm{\tau},H_{1:t})$ in the task space.
@@ D.1 Neural Modules to Parameterize Distributions
Here, we detail the neural modules to parameterize the distributions of interest. For the approximate posterior $q_{\bm{\phi}}(\bm{z}_{t}|H_{t})$ and conditional prior $p(\bm{z}_{t}|H_{1:t-1})$, the inputs of the module are a set of task risk pairs. The neural module requires the permutation invariance w.r.t. the order of the data points in the set $H_{t}$ or $H_{1:t-1}$ in Definition 3. Hence, we adopt the DeepSet style neural network 56 to process the collected $H_{t}$ or $H_{1:t-1}$.
Definition 3 (Permutation Invariant Function)
$g:[1,2,\dots,n]\mapsto[g_{1},g_{2},\dots,g_{n}].$ (23)
@@
Connections with the NP Family. Regarding the task risk functional posterior inference module, this work has a close connection with the NP family 81, 55, 91, 92, 93, 94, 79. Both handle the set data points in probabilistic inference.
Also, note that this work is significantly distinguished from the NP in terms of research motivations and learning setups. The NP mostly copes with the stationary stochastic process and approximates the underlying functional prior with the help of a context-based latent variable. Our setup considers learning in the non-stationary case, where histories arrive sequentially and depend on the updated machine learner. Hence, we adopt the streaming VI 40 and capture the task risk value in a posterior inference manner.
@@ D.2 Formulation of ELBO & Stochastic Gradient Estimates
Then, we can rewrite the ELBO with the help of reparameterization trick 25 in Eq. (27).
$\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\Big{[}\ln p_{\bm{\psi}}(H_{t}|\bm{z}_{t})\Big{]}-D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel p(\bm{z}_{t}|H_{1:t-1})\Big{]}$ (27a)
@@ D.3 Theoretical Guarantee for Task Difficulties’ Scoring with Posterior Inference
Given the Assumption 3 and the Chernoff bound 95, we can have the concentration inequality as:
@@
The gradient descent as fast adaptation:
@@ Appendix E Prompt-based Few-shot Image Classification
We adopt the standard few-shot image classification setting 28, 85, where tasks are constructed using the N-way K-shot paradigm for both meta-training and meta-testing. Each task comprises support and query sets. The support set contains K examples for each of the N classes, while the query set includes 15 examples per class. During meta-training, labels for both support and query data are accessible to the adaptive machine learner. During meta-testing, the query dataset’s labels are to be predicted given the labeled support dataset. The class categories of task datasets in the meta-training and meta-testing do not overlap. In experiments, we specifically consider a 5-way 1-shot image classificationconfiguration. During meta-training, we set the task batch for ERM and DRM as $\mathcal{B}=4$ (For implementation simplicity, the data loader samples $8$ tasks and then randomly keeps half without ranking to optimize). The task batch for DRM is $\mathcal{B}=8$ before the filtering operation; DRM filters half to optimization. Similarly, that for MPTS is $\mathcal{B}=8$ and only $4$ tasks are screened to optimize.
To enable few-shot learning by prompt-tuning, we integrate the multimodal prompt learning methods MaPLe 44 and prototypical network (ProtoNet) 85. MaPLe operates on the CLIP model 1, capturing multimodal prompts to refine both visual and textual feature representations with frozen CLIP’s parameters. These refined textual features serve as classifiers for predicting refined image features. In parallel, ProtoNet is utilized to derive class-specific visual embeddings from the support set, which assist in distinguishing query samples.

### v5

@@ Appendix B Research Background
B.1 Adaptation Learning for Cross-Task Generalization
Learning from zero-shot or few-shot examples has been identified as a crucial adaptation capability of the machine learner nowadays 68, 69. In SFT, this work treats the individual example as each task to meet MPTS setup. As SFT techniques have been widely discussed in the field 56, we skip this part in the background introduction.
@@ B.1 Adaptation Learning for Cross-Task Generalization
This assesses the machine learner’s generalization capability when directly deploying in unseen scenarios without the help of a support dataset. Such a cross-task generalization is commonly studied in computer vision 70, and the core of the relevant methods is effective semantic representation either from embedding-based methods 71, 72, 73 or generative-based methods 74, 75. In the era of the foundation models, the pretraining mechanism between multimodality also sometimes empowers the machine learner, such as CLIP 1, with zero-shot capability. When it comes to sequential decision-making, a commonly seen method is DR 55, 35, which places a distribution over environments for the agent to interact.
@@
This examines the machine learner’s capability of resolving unseen tasks from some annotated examples as hints. Meta-learning, as the typical learning paradigm, has gained popularity over the past decade. It achieves few-shot adaptation by leveraging past experience and distilling knowledge to unseen but similar scenarios in a few-shot way 76. In brief, we categorize commonly seen methods into context-based, optimization-based, geometric-based, and others. (i) Formulated in an encoder-decoder structure, the context-based method resembles variational autoencoders and encodes the few-shot information into latent variables or embeddings. Typical ones are neural process families 50, 77, 78, 79, which aim to constitute exchangeable deep stochastic processes with neural networks. (ii) The optimization-based methods, with their versatile nature and ability to enable cross-task skill transfer, have piqued the interest and engagement of researchers in the field. For example, MAML 29, 80, 81, 82 reduces meta-learning to a bi-level optimization in the parameter space, and its extensions have been widely investigated in the field. (iii) The deep metric-based methods 83, 84 attempt to embed tasks into the latent space and are more suitable for few-shot image classification tasks. Besides, there are other families, such as hyper-networks 85, 86, recurrent meta-learning 7, etc.
@@ B.2 Dataset Curation and Task-Level Robustness
Task Curation in Robust Adaptation Learning Pipelines.
Recent works 15, 16 demonstrate the effectiveness of challenging task prioritization over uniform sampling in improving cross-task generalization and adaptation robustness, particularly when the learning dataset is sufficiently large. Many methods 39, 15, 16, 14, 13, 17, 18 adopt an Evaluate-Rank-Filter step for iterative model updates, introducing a batch filtering ratio $\hat{\alpha}=1-\frac{\mathcal{B}}{\hat{\mathcal{B}}}\in[0,1)$ to quantify the fraction of discarded tasks in a sample batch. This prioritization of "difficult" tasks aligns with minimizing $\text{CVaR}_{\alpha}$ 30, a robustness metric for tail-case performance. Alternatively, other methods 41, 42, 43 focus on constructing uncertainty sets and reweighting tasks within the batch to achieve robust adaptation. Additionally, coreset methods 61, 62, 63 aim to select a small subset of tasks that effectively represent the utility of the full dataset, often through gradient approximation in optimization. These approaches address a subproblem of data efficiency, with the acquisition strategy in MPTS serving as an episodic coreset selection mechanism tailored for robustness.
@@
The $\text{CVaR}_{\alpha}$ or expected shortfall 30 is a statistical measure to assess the proportional worst-case performance of some models at certain levels. This is widely adopted in risk-averse applications. As implied in Definition 1, $\text{CVaR}_{\alpha}$ describes the expected risk under the normalized $(1-\alpha)$ proportional tail risk task distribution, and this work specifies the distribution in the task space. Meanwhile, the normalized tail task distribution $p_{\alpha}(\tau;\bm{\theta})$ can be viewed as a shifted result from the initial task distribution $p(\tau)$; hence, such a measure provides robustness quantification in the presence of task distribution shifts 14, 40, 17.
Another indicator to evaluate the machine learner’s robustness is the performance in OOD tasks. This refers to the case when the training and the testing task distributions are different. Particularly, in DR and prompt-tuning scenarios, we also use the OOD tasks that never appear in the training task distribution to test the trained policy, and this setup corresponds to domain generalization, a type of substantial distribution shift 8.
B.3 Risk Minimization Principles as Baselines
Expected/Empirical Risk Minimization (ERM).
With the fixed $p(\tau)$, the principle follows the statistical learning theory 38 and minimizes the expectation of adaptation risk over the task space. As a result, we can have:
It draws batches with a random task sampler to optimize iteratively.
Distributionally Robust Risk Minimization (DRM) 39, 16, 17, 14, 40.
We retain the notation of task robust work 14, which terms the tail task risk minimization as DRM. It aims to improve the robustness of adaptation to the tail tasks over iteration. No explicit form exists as the tail task distribution is $\bm{\theta}$-dependent. The optimization objective is derived as the $\text{CVaR}_{\alpha}(\bm{\theta})$ 30:
where we write $p_{\alpha}(\tau;\bm{\theta})$ to express the $(1-\alpha)$ proportional worst case for easier formulation. In other words, $\mathbb{E}_{p_{\alpha}(\tau;\bm{\theta})}\Big{[}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\Big{]}$ also relates to the task distribution with constraints. Also note that when $\alpha$ approaches $1$, the problem degenerates to the worst-case risk minimization.
This work retains the setup in work 14 and picks up the $-\mathcal{B}$ in optimization, which corresponds to sample average Monte Carlo of $\text{CVaR}_{\alpha}$. This implies that the actual task batch to evaluate is $\frac{\mathcal{B}}{1-\alpha}$. And for fair comparison with MPTS and light computations, we retain the Monte Carlo estimator for the risk quantile in implementation. To ensure stable training, in all benchmarks, we keep the actual task batch ^$\hat{\mathcal{B}}=2\mathcal{B}$ to evaluate and discard the easiest half before the machine learner’s optimization.
Group Distributionally Robust Risk Minimization (GDRM) 41.
This can be interpreted as a min-max optimization problem. Such a principle 41 effectively improves robustness in distribution shifts and has shown positive effects on training foundation models 42, 43. It constructs a collection of uncertainty sets over tasks and results in the optimization objective as follows:
GDRM increases the machine learner’s robustness by assigning more probability mass to worst cases in a reweighted manner. That means in each iteration with the best selected $p_{\hat{g}}(\tau)$, the optimization problem is reduced to
@@ B.3 Risk Minimization Principles as Baselines
Given a fixed number of tasks, GDRM will heuristically or dynamically group them into clusters and then perform a reweighting mechanism according to the evaluated risk. In task episodic learning, there is no task grouping operation as the task batch is reset after each iteration. And the default computation of task-specific weights is $\omega(\tau_{i})=\frac{\exp(\eta\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}))}{\sum_{b=1}^{\mathcal{B}}\exp(\eta\ell(\mathcal{D}_{\tau_{b}}^{Q},\mathcal{D}_{\tau_{b}}^{S};\bm{\theta}))}$, where $\eta$ is the temperature parameter and $\{\bm{\tau}_{b}\}_{b=1}^{\mathcal{B}}$ is the identifier of the task batch. The implementation detail can be found in https://github.com/kohpangwei/group_DRO.
As revealed in works 87, 88, 14, the heuristic operation as the Evaluate-Rank-Filter or reweighting mechanism in GDRM is widely adopted for approximate optimization. For example, in task robust meta-learning scenarios, the prerequisite step in DR-MAML 14 is to execute gradient updates in the inner loop for all candidate tasks and then screen the tail task subset to meta-optimize according to the evaluation results.
Appendix C Task Construction and Identifiers
@@ C.1 Tasks with Explicit Identifiers
In this setup 29, meta learners aim at quickly adapting the model to an unseen function $f(x)=a\sin(x-b)$ with the help of $K$ data points randomly sampled from the function. This case treats the amplitude and phase variables $(a,b)$ as the task identifier to configure the task. And the task distribution is induced by the uniform distribution over the task identifier.
@@
DR is a promising paradigm to achieve zero-shot adaptation in unseen scenarios, which is widely adopted in robotics 35 and computer vision 55. The basic idea is to train the machine learner in a distribution over environments and then directly apply the learned model to new ones.
@@ C.2 Tasks with Implicit Identifiers
Nevertheless, in several scenarios, it is intractable to access the explicit task identifier. For example, in few-shot image classification, the task information is just the coupled support and query dataset $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$. Similarly, in SFT for LLMs, the task can be in the form of the QA pair $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{Q}$. There is no explicit representation method, such as $\bm{\tau}$, for these tasks, which brings difficulty in building up the risk learner. Retaining the prior notation, the episodic task batch can be written as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$, where $\bm{\tau}$ of our interest is unobservable. Some experiments in this work, such as few-shot image classification and SFT, encounter such circumstance.
@@ Appendix D Auto-Encoding Adaptation Risk through Streaming VI
Note that the basis of MPTS is to establish the bridge between the task identifier and the adaptation risk value over the course of the machine learner’s optimization. In other words, we are seeking a lightweight stochastic risk function in Definition 3 to approximate the posterior $p(\ell|\bm{\tau},H_{1:t})$ in the task space.
@@ D.1 Neural Modules to Parameterize Distributions
Here, we detail the neural modules to parameterize the distributions of interest. For the approximate posterior $q_{\bm{\phi}}(\bm{z}_{t}|H_{t})$ and conditional prior $p(\bm{z}_{t}|H_{1:t-1})$, the inputs of the module are a set of task risk pairs. The neural module requires the permutation invariance w.r.t. the order of the data points in the set $H_{t}$ or $H_{1:t-1}$ in Definition 2. Hence, we adopt the DeepSet style neural network 51 to process the collected $H_{t}$ or $H_{1:t-1}$.
@@
Regarding the task risk functional posterior inference module, this work has a close connection with the NP family 79, 50, 89, 90, 91, 92, 77. Both handle the set data points in probabilistic inference.
@@ D.2 Formulation of ELBO & Stochastic Gradient Estimates
Then, we can rewrite the ELBO with the help of reparameterization trick 23 in Eq. (27).
$\hat{\mathcal{G}}_{\text{ELBO}}(\bm{\psi},\bm{\phi})=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\Big{[}\ln p_{\bm{\psi}}(H_{t}|\bm{z}_{t})\Big{]}-D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel p(\bm{z}_{t}|H_{1:t-1})\Big{]}$ (27a)
@@ D.3 Theoretical Guarantee for Task Difficulties’ Scoring with Posterior Inference
Given the Assumption 3 and the Chernoff bound 93, we can have the concentration inequality as:
@@
The gradient descent as fast adaptation is denoted by:
@@ Appendix E Prompt-based Few-shot Image Classification
We adopt the standard few-shot image classification setting 29, 83, where tasks are constructed using the N-way K-shot paradigm for both meta-training and meta-testing. Each task comprises support and query sets. The support set contains K examples for each of the N classes, while the query set includes 15 examples per class. During meta-training, labels for both support and query data are accessible to the adaptive machine learner. During meta-testing, the query dataset’s labels are to be predicted given the labeled support dataset. The class categories of task datasets in the meta-training and meta-testing do not overlap. In experiments, we specifically consider a 5-way 1-shot image classificationconfiguration. During meta-training, we set the task batch for ERM and DRM as $\mathcal{B}=4$ (For implementation simplicity, the data loader samples $8$ tasks and then randomly keeps half without ranking to optimize). The task batch for DRM is $\mathcal{B}=8$ before the filtering operation; DRM filters half to optimization. Similarly, that for MPTS is $\mathcal{B}=8$ and only $4$ tasks are screened to optimize.
To enable few-shot learning by prompt-tuning, we integrate the multimodal prompt learning methods MaPLe 36 and prototypical network (ProtoNet) 83. MaPLe operates on the CLIP model 1, capturing multimodal prompts to refine both visual and textual feature representations with frozen CLIP’s parameters. These refined textual features serve as classifiers for predicting refined image features. In parallel, ProtoNet is utilized to derive class-specific visual embeddings from the support set, which assist in distinguishing query samples.